OpenStudy (anonymous):

Express y in terms of x log y - log x = 5 log 3

OpenStudy (anonymous):

log(a) + log(b) = log(ab) log(a) - log(b) = log(a/b) log(a^b) = b*log(a) using second and third rule: \[\log(y/x) = \log(3^5)\]\[\frac{ y }{ x } = 3^5\]\[y = 243x\]